• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Carlos

    Log entry time 23:57:48 on November14,2004

    Entry number 136679

    keyword=Swing shift summary

    Shift crew:
    ===========
    C. Munoz Camacho
    C. Ferdi
    P. Ulmer

    Shift summary:
    ==============

    Pretty smooth shift. Steady beam during the whole shift. Data taking went without big issues. DAQ was rock solid. Only a couple of IOC reboots and one Compton reboot per run...

    ---
    16:00 Shift started with production data on LD2.

    19:19 Compton needed coda reboot (see halog entry 136660).

    20:00 DVCS IOC freezes. Rebooted and working.

    21:03 DVCS IOC froze again. Rebooted.

    21:44 Bleedthrough measurement: 2.90% (ELOG entry 1244871).

    21:55 Compton crashed again. Coda rebooted. (see halog entry 136669).

    22:13 Asked MCC to tune beam vertical position to 1.40 in BPM4B-Y (instead of 1.19).

    22:23 BPM positions corrected. It seems that the adjustements of beam position make the beam trip much more than usual (4 times in 10min of adjustement). This happened yesterday also when I asked for that.

    22:33 LD2 TWOARM run

    22:39 Back to LD2 production runs.

    22:50 Momentarily high dead time according to datamon (see